is_pk is min isl_pk i min Fig. 2

Question 1 For the flyback converter shown in Fig. 1, Vin is 30 V, Vout is 8 V, the output power is 30 w, 30 turns are fitted to winding Ww1, 15 turns are fitted to winding W2, the inductance of W1, L1, is 50 uH and the switching frequency is 200 kHz. The flyback converter operates in continuous conduction mode. Calculate the time for which S1 is switched on a. b. Calculate Aisı during the time S1 is switched on c. Calculate is1_min, isı_pk, is2_min and is2_pk, as shown in Fig. 2 d. Calculate the minimum blocking voltage rating needed for S2, assuming no transient voltages are generated when it switches on and off. S2 Vin Vout Load W2 C Si Fig. 1
is_pk is min isl_pk i min Fig. 2
Answer #1

+ly BACK Conuerter i. AN S2 HhoS 301S Load Yout -bvolt utyra tio Let D be Vin Ton doo KHz Ton 3T Ls. NN2 in for fyback Convesmnzياin t T isimin- T Nin t 1s2-p Ismin when Switch s, is on then Szis o across it- Vs,- 2-No the veverse -Vo -N Vin Kalue o(39 3,6 1 -2 8-8D- 1SD tia D= 8 duty 3 (4) Time S, s on To n Ton =1.73913 s T= 200x 103 during VLw Vin: Lun at on diu Ais - (IL = 30 IL 23 Q-B 75 A So wp Is_pK = IL Is, Pr o 43478 = 3.39 67 A <. 87S Iwmin Is,_min = IL-Aiw, s 875- 1-o4378 Ts, _min a dSo N2 = 6.793 A isspe So Зо 1 ismin sin = N_(\Lw,)min = is2_m Nou N2 Nz. 4.7062A 30 x.3 S31 Sa ainaffstate mode i.e blacking